OpenStudy (anonymous):

What is 4x + 2y = -20 written in slope-intercept form? A. y = 2x - 20 B. y = -4x + 20 C. y = 20x - 4 D. y = -2x - 10

OpenStudy (anonymous):

rearrange 4x+2y=-20 into the form of y=mx+c :)

OpenStudy (anonymous):

2y=-4x-20 y=-2x-10

OpenStudy (anonymous):

D

OpenStudy (anonymous):

d